OpenStudy (anonymous):

okay over thinking 3 1/2 + 5/6 =

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\rightarrow 3 + {1 \over 2} + {5 \over 6} }\) \(\Large \color{MidnightBlue}{\rightarrow 3 + {3 \over 6} + {5 \over 6} }\) \(\Large \color{MidnightBlue}{\rightarrow 3 + {8 \over 6} }\) \(\Large \color{MidnightBlue}{\rightarrow 3 + 1{1 \over 2} }\) \(\Large \color{MidnightBlue}{\rightarrow 4{1 \over 2} }\)

OpenStudy (anonymous):

3 1/2 + 5/6 = 7/2 + 5/6( find common denominator of 2 and 6 in this case it's 6) =(21+5)/6 = 26/6 =4 1/3

OpenStudy (lgbasallote):

@ParthKohli \[\large \frac{8}{6} = 1 \frac{1}{3}\] ^_^

Parth (parthkohli):

Oh hell noo